Assume that we want to construct a confidence interval. Do one of the​ following, as​ appropriate: (a) find the critical value t Subscript alpha divided by 2​, ​(b) find the critical value z Subscript alpha divided by 2​, or​ (c) state that neither the normal distribution nor the t distribution applies. Here are summary statistics for randomly selected weights of newborn​ girls: nequals181​, x overbarequals33.8 ​hg, sequals6.2 hg. The confidence level is 99​%.
Answer #1

Date Solution- Given that, x=33.8 s= 6-2 n=181 df-n-1 =181-1=180 At aq olo confidence level the tis 2-1-99% 221-0.99 = 0, 0]